# Jean DELHAYE

### v2

@@ 3.1. Defining the quantum unitary Brownian motion
This further strengthens the justification for labeling functionals of the form (3.2) as Brownian motions.
We are not asserting that every pair $(\alpha,\beta)$ satisfying $\alpha\geq\beta/N\geq 0$ necessarily defines a central generating functional through Formula (3.2), although this is something we anticipate to be true. Instead, we are stating that all generating functionals of this form should be regarded as Brownian motions. The existence of such examples is guaranteed by the Brownian motion on $U_{N}$. Similarly, processes where $\alpha\geq 0$ and $\beta=0$ can be constructed by composing a Brownian motion on $O_{N}^{+}$ (a generating functional described by Formula (3.1) with $\nu=0$) with the quotient map $\mathcal{O}(U_{N}^{+})\to\mathcal{O}(O_{N}^{+})$ and the conditional expectation $\mathbb{E}:\mathcal{O}(U_{N}^{+})\to\mathcal{O}(U_{N}^{+})_{0}$. From these two types of examples, many more can be generated, as the set of all generating functionals forms a convex cone.
Definition 3.2.
Any central generating functional on $U_{N}^{+}$ of the form
$L_{\alpha,\beta}:\chi_{w}\longmapsto-{\ell(w)}N^{\ell(w)-1}\alpha-\Big{(}{(p(w)-q(w))^{2}-\ell(w)}\Big{)}N^{\ell(w)-2}\beta,\quad w\hbox{ word on }\ \{\lozenge,\blacklozenge\},$
for some $\alpha,\beta\geq 0$ is called Brownian motion of parameter $(\alpha,\beta)$. We will also call Brownian motion of parameter $(\alpha,\beta)$, the associated Lévy process.
@@ 3.2. Computing the values of the Brownian motion
Let $L=L_{\alpha,\beta}$ be a Brownian motion on $U_{N}^{+}$. The goal of this subsection is to compute the values of $L$ on all characters of irreducible representations. Note that $L$ coincides with a Gaussian generating functional on the central algebra. Therefore, we may apply Lemma A.3 to $L$ on elements of the central algebra. Our first goal is to understand its values on elementary elements and products of two elementary elements. Specifically, we aim to determine the quantities of the form
@@
(3.3) $\Phi_{m}^{\epsilon_{1}}=-{m}N^{m-1}\alpha-({\mathfrak{p}_{m}-m})N^{m-2}\beta,$
@@
(3.4) $\Phi_{mn}^{\epsilon_{1}\eta_{1}}=-({m+n})N^{m+n-1}\alpha-\Big{(}{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}-(m+n)}\Big{)}N^{m+n-2}\beta$
@@
$\mathfrak{p}_{m}=\left\{\begin{array}[]{ll}0&\hbox{if }m\hbox{ is even}\\ 1&\hbox{else}\end{array}\right.\quad\hbox{\&}\quad\mathfrak{p}_{mn}=\left\{\begin{array}[]{ll}0&\hbox{if }m\hbox{ or }n\hbox{ is even}\\ 2&\hbox{else}\end{array}\right.$
@@
Let $w$ be the word such that $z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}=\chi_{w}$. It is not hard to see that $w$ is a word of alternating letters of length $m$. Particularly, $\ell(w)=m$ and $d(w)=\mathfrak{p}_{m}=d(w)^{2}$. Now, it is clear that Formula (3.3) follows from Equation (3.2).
@@
We can now understand the Brownian motion on some irreducible characters.
@@
(3.5) $\Psi_{m}^{\epsilon_{1}}:=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}})=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N^{2}}P_{m}(N),$
@@
(3.6) $\begin{split}\Psi_{mn}^{\epsilon_{1}\eta_{1}}:&=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi_{n}z^{[\eta_{2}]_{+}})\\ &=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N^{2}}(P_{m}P_{n})(N)\end{split}$
@@
$=\sum_{j}a_{j}^{m}\Big{(}-jN^{j-1}\alpha-(\mathfrak{p}_{j}-j)N^{j-2}\beta\Big{)}$
$=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\sum_{j}a_{j}^{m}\mathfrak{p}_{m}N^{j-2}\beta$
$=-\left(\alpha-\frac{\beta}{N}\right)P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N^{2}}P_{m}(N).$
@@
$=\sum_{i,j}a_{i}^{m}a_{j}^{n}\left(-(i+j)N^{i+j-1}\alpha-\Big{(}{(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij}-(i+j)}\Big{)}N^{i+j-2}\beta\right)$
$=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\sum_{i,j}a_{i}^{m}a_{j}^{n}\Big{(}{(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij}}\Big{)}N^{i+j-2}\beta$
$=-\left(\alpha-\frac{\beta}{N}\right)(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N^{2}}(P_{m}P_{n})(N).$
Before stating this section’s final theorem, we need one last ingredient. We first need to define the parity entanglement of a tuple $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, to do so let us fix some additional notations.
@@
$\mathfrak{e}_{\mathbf{n}}:=\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}.$
@@
Let $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$ and recursively define $\epsilon_{i+1}:=(-1)^{n_{i}+1}\epsilon_{i}$, then
$\sum_{1\leq i<j\leq p}\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}=2\mathfrak{e}_{\mathbf{n}}.$
@@
We may assume that $\mathfrak{p}_{\mathbf{n}}\geq 2$ since both terms are null otherwise. First note that for any $1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}$:
$(-1)^{n_{k_{i}}+...+n_{k_{j-1}}}=(-1)^{j-i}.$
This is because the amount of odd numbers in $\{n_{k_{s}}:i\leq s<j\}$ is exactly $j-i$. This further implies
$\epsilon_{k_{i}}\epsilon_{k_{j}}$ $=(-1)^{n_{k_{i}}+...+n_{k_{j-1}}+k_{j}-k_{i}}\epsilon_{k_{i}}^{2}=(-1)^{n_{k_{i}}+...+n_{k_{j}-1}+k_{j}-k_{i}}=(-1)^{k_{j}+j-(k_{i}+i)}.$
@@
$=2\sum_{1\leq j<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}$
$=2\mathfrak{e}_{\mathbf{n}}.$
From the second to third line, we have used the fact that $12\delta_{xy}=xy+$ 1 when $x,y$ are elements in {±$\{\pm 1\}$. Thus, the equality holds. ∎
@@
Let $\mathbf{n}=(n_{1},...,n_{p})\in\mathbb{N}^{*p}$, $\epsilon\in\{\pm 1\}$, then
$L(\chi^{\epsilon}_{\mathbf{n}})=-\left(\alpha-\frac{\beta}{N}\right)P^{\prime}_{\mathbf{n}}(N)-\beta\frac{\mathfrak{p}_{\mathbf{n}}+2\mathfrak{e}_{\mathbf{n}}}{N^{2}}P_{\mathbf{n}}(N),$
where $P_{\mathbf{n}}=P_{n_{1}}...P_{n_{p}}$.
@@
$L(\chi_{\mathbf{n}}^{\epsilon})$ $=L(z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{[\epsilon_{2}]_{+}}...z^{[\epsilon_{p}]_{-}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}})$
@@
$\,\,\,\,\,\,\,\,\,\,\,\,\,-(p-1)\sum_{j=1}^{p}\left(-\left(\alpha-\frac{\beta}{N}\right)\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}-\beta\frac{\mathfrak{p}_{n_{j}}}{N^{2}}\right)d_{\mathbf{n}}$
$=-\left(\alpha-\frac{\beta}{N}\right)\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\mathfrak{p}_{n_{i}}+\mathfrak{p}_{n_{j}}-2\mathfrak{p}_{n_{i}}\mathfrak{p}_{n_{j}}+\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}^{2}_{n_{i}n_{j}}}{N^{2}}d_{\mathbf{n}}$
$\,\,\,\,\,\,\,\,\,\,\,\,\,+\beta(p-1)\sum_{j=1}^{p}\frac{\mathfrak{p}_{n_{j}}}{N^{2}}d_{\mathbf{n}}$
$=-\left(\alpha-\frac{\beta}{N}\right)\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\frac{\mathfrak{p}_{\mathbf{n}}}{N^{2}}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}}{N^{2}}d_{\mathbf{n}}$

### v3

@@ 3.1. Defining the quantum unitary Brownian motion
We gather the results of Proposition 3.2, Proposition A.4 together with the subsequent comments, into a single theorem to highlight our definition of Brownian motion on $U_{N}^{+}$.
Theorem 3.3.
Every pair $(\alpha,\beta)\in\mathbb{R}^{2}$ defines a central generating functional on $U_{N}^{+}$ through the formula
(3.2) $L_{\alpha\beta}:\chi_{w}\longmapsto-\Big(\ell(w)\alpha+\big(p(w)-q(w)\big)^{2}\beta\Big)N^{\ell(w)-1}.$
Moreover:
  1. (i)
    They exactly describe (real) centralized Gaussian generating functionals.
  2. (ii)
    The centralized pulled-back Brownian motion from the classical group $U_{N}$ corresponds to such a generating functional.
We call such generating functionals (along with the associated Lévy processes), Brownian motions.
@@ 3.2. Computing the values of the Brownian motion
Let $L=L_{\alpha\beta}$ be a Brownian motion on $U_{N}^{+}$ as defined in Theorem 3.3. The goal of this subsection is to compute the values of $L$ on all characters of irreducible representations. By definition, $L$ coincides with a Gaussian generating functional on the central algebra. Therefore, we may apply Lemma A.3 to $L$ on elements of the central algebra. Our first goal is to understand its values on elementary elements and products of two elementary elements. Specifically, we aim to determine the quantities of the form
@@
(3.3) $\Phi_{m}^{\epsilon_{1}}=-\big(m\alpha+\mathfrak{p}_{m}\beta\big)N^{m-1},$
@@
(3.4) $\Phi_{mn}^{\epsilon_{1}\eta_{1}}=-\Big(({m+n})\alpha+\big((\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}\big)\beta\Big)N^{m+n-1}$
@@
[…]
@@
For any word $w$ in the letters $\{\lozenge,\blacklozenge\}$, we set d(w):=|p(w)q(w)|d(w):=|p(w)-q(w)|.
Let $w$ be the word such that $z^{[\epsilon_{1}]_{-}}\chi^{m}_{1}z^{[\epsilon_{2}]_{+}}=\chi_{w}$. Note that $w$ is a word of alternating letters of length $m$. Particularly, $\ell(w)=m$ and $d(w)=\mathfrak{p}_{m}=d(w)^{2}$. Now, it is clear that Formula (3.3) follows from Equation (3.2).
@@
We now compute the Brownian motion on some irreducible characters.
@@
(3.5) $\Psi_{m}^{\epsilon_{1}}:=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}})=-\alpha P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N}P_{m}(N),$
@@
(3.6) $\begin{split}\Psi_{mn}^{\epsilon_{1}\eta_{1}}:&=L(z^{[\epsilon_{1}]_{-}}\chi_{m}z^{[\epsilon_{2}]_{+}}\cdot z^{[\eta_{1}]_{-}}\chi_{n}z^{[\eta_{2}]_{+}})\\ &=-\alpha(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N}(P_{m}P_{n})(N)\end{split}$
@@
$=-\sum_{j}a_{j}^{m}\Big(j\alpha+\mathfrak{p}_{j}\beta\Big)N^{j-1}$
$=-\alpha P_{m}^{\prime}(N)-\sum_{j}a_{j}^{m}\mathfrak{p}_{m}N^{j-1}\beta$
$=-\alpha P_{m}^{\prime}(N)-\beta\frac{\mathfrak{p}_{m}}{N}P_{m}(N).$
@@
$=-\sum_{i,j}a_{i}^{m}a_{j}^{n}\left((i+j)\alpha+\Big({(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij})}\Big)\beta\right)N^{i+j-1}$
$=-\alpha(P_{m}P_{n})^{\prime}(N)-\sum_{i,j}a_{i}^{m}a_{j}^{n}\Big({(\mathfrak{p}_{i}-\mathfrak{p}_{j})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}^{2}_{ij}}\Big)\beta N^{i+j-1}$
$=-\alpha(P_{m}P_{n})^{\prime}(N)-\beta\frac{(\mathfrak{p}_{m}-\mathfrak{p}_{n})^{2}+\delta_{\epsilon_{1}\eta_{1}}\mathfrak{p}_{mn}^{2}}{N}(P_{m}P_{n})(N).$
Before stating this subsection’s final theorem, we need one last ingredient. We first need to define the parity entanglement of a tuple $\mathbf{n}=(n_{1},\cdots,n_{p})\in\mathbb{N}^{*p}$, to do so let us fix some additional notations.
@@
$\mathfrak{e}_{\mathbf{n}}:=\sum_{1\leq i,j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}.$
@@
Let $\mathbf{n}=(n_{1},\cdots,n_{p})\in\mathbb{N}^{*p}$, $\epsilon=\epsilon_{1}\in\{\pm 1\}$ and recursively define $\epsilon_{i+1}:=(-1)^{n_{i}+1}\epsilon_{i}$, then
$\mathfrak{p}_{\mathbf{n}}+\sum_{1\leq i<j\leq p}\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}=\mathfrak{e}_{\mathbf{n}}.$
@@
Let $k_{1},\cdots,k_{\mathfrak{p}_{\mathbf{n}}}$ denote the indices of the odd entries of $\mathbf{n}$.
If $\mathfrak{p}_{\mathbf{n}}=0$, both terms of the equality are null, if $\mathfrak{p}_{\mathbf{n}}=1$, both terms equal $1$. We assume that $\mathfrak{p}_{\mathbf{n}}\geq 2$. First note that for any $1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}$:
$(-1)^{n_{k_{i}}+\cdots+n_{k_{j-1}}}=(-1)^{j-i}.$
This is because the $n_{k_{s}}$’s are odd. This further implies
$\epsilon_{k_{i}}\epsilon_{k_{j}}$ $=(-1)^{n_{k_{i}}+\cdots+n_{k_{j-1}}+k_{j}-k_{i}}\epsilon_{k_{i}}^{2}=(-1)^{n_{k_{i}}+\cdots+n_{k_{j}-1}+k_{j}-k_{i}}=(-1)^{k_{j}+j-(k_{i}+i)}.$
@@
$=2\sum_{1\leq i<j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}$
$=\sum_{1\leq i,j\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{j}+j-(k_{i}+i)}-\sum_{1\leq i\leq\mathfrak{p}_{\mathbf{n}}}(-1)^{k_{i}+i-(k_{i}+i)}$
$=\mathfrak{e}_{\mathbf{n}}-\mathfrak{p}_{\mathbf{n}}.$
From the second to third line, we have used the fact that $2\delta_{xy}=xy+1$ when $x,y$ are elements in $\{\pm 1\}$. Thus, the equality holds. ∎
@@
Let $\mathbf{n}=(n_{1},\cdots,n_{p})\in\mathbb{N}^{*p}$, $\epsilon\in\{\pm 1\}$, then
$L(\chi^{\epsilon}_{\mathbf{n}})=-\alpha P^{\prime}_{\mathbf{n}}(N)-\beta\frac{\mathfrak{e}_{\mathbf{n}}}{N}P_{\mathbf{n}}(N),$
where we have written $P_{\mathbf{n}}=P_{n_{1}}\cdots P_{n_{p}}$.
@@
$L(\chi_{\mathbf{n}}^{\epsilon})$ $=L(z^{[\epsilon_{1}]_{-}}\chi_{n_{1}}z^{[\epsilon_{2}]_{+}}\cdots z^{[\epsilon_{p}]_{-}}\chi_{n_{p}}z^{[\epsilon_{p+1}]_{+}})$
@@
$\,\,\,\,\,\,\,\,\,\,\,\,\,-(p-1)\sum_{j=1}^{p}\left(-\alpha\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}-\beta\frac{\mathfrak{p}_{n_{j}}}{N}\right)d_{\mathbf{n}}$
$=-\alpha\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\mathfrak{p}_{n_{i}}+\mathfrak{p}_{n_{j}}-2\mathfrak{p}_{n_{i}}\mathfrak{p}_{n_{j}}+\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}^{2}_{n_{i}n_{j}}}{N}d_{\mathbf{n}}$
$\,\,\,\,\,\,\,\,\,\,\,\,\,+\beta(p-1)\sum_{j=1}^{p}\frac{\mathfrak{p}_{n_{j}}}{N}d_{\mathbf{n}}$
$=-\alpha\sum_{j=1}^{p}\frac{P_{n_{j}}^{\prime}(N)}{P_{n_{j}}(N)}d_{\mathbf{n}}-\beta\frac{\mathfrak{p}_{\mathbf{n}}}{N}d_{\mathbf{n}}-\beta\sum_{1\leq i<j\leq p}\frac{\delta_{\epsilon_{i}\epsilon_{j}}\mathfrak{p}_{n_{i}n_{j}}^{2}-\mathfrak{p}_{n_{i}n_{j}}}{N}d_{\mathbf{n}}$